If 50.0 mL of 0.125 mol/L sodium phosphate solution is reacted with 50.0 mL of 0.125 mol/L magnesium nitrate solution, what is the predicted mass of precipitate in g?
1
Expert's answer
2020-06-15T14:43:11-0400

  Solution.

 0.00625               0,00625                           y

2Na3PO4 + 3Mg(NO3)2 → Mg3(PO4)2↓ + 6NaNO3

      2                               3                               1

V(Na3PO4) = 50,0 mL = 0,05 L;  

V(Mg(NO3)2) = 50,0 mL = 0,05 L;

n(Na3PO4) = c*V = 0,125*0,05 = 0,00625 mol;

n(Mg(NO3)2) = c*V = 0,125*0,05 = 0,00625 mol;

Let n(Mg(NO3)2) is equal to x, therefore x = 0,00625*3/2 = 0,009375 mol.

We have only 0,00625 mol Mg(NO3)2;

Mg(NO3)2 – in lack.

n(Mg3(PO4)2) = y = 0,00625*1/3 = 0,002083 mol;

m(Mg3(PO4)2) = 0,002083*263 = 0,548 g.

 Answer: 0,548 g Mg3(PO4)2.